A car is initially at rest on a straight road. The histogram below shows the cars acceleration along that road as a function of time. 1 -3 .1234567ธ9101112 14 Tine (s) Calculate the speed of the car at t3s. Your receipt no. is 158-5693Previous Tries 6.00 You are correct. Calculate the distance traveled during the first 5 s. 24.0 m ubAnwer Incorrect. Tries 19/20 Previous Tries Calculate the distance traveled from t-10 s to t=14 s 36 m The car does not accelerate during that time interval and therefore its speed is constant. Submit Answer Incorrect. Tries 6/20 Previous Tries Calculate the cars average speed fromt 6 s to t-9 s. The average speed is the distance traveled divided by the time elapsed. Try using this graph to create a velocity vs. time graph, where the slopes of the new graph are given by the 1.00 m/s accelerations on this graph Submit Answer Incorrect. Tries 2/20 Previous Tries
Answer #1

As you know speed

v=x/t means if t=1 s then v=x

acceleration a=v/t and v= at

Uine euntion ot the mshion x,1x 13.sm hen of the fstnho Vectanale first S see. Speed atssee is nls Ginal uend equation tf the motion 32m ave sage speed 十ravel distance G-(ㅇㄱ is toled dis temce tott fine 8-7 1 from O 2 g+16.stt‘ー 23.s. D.> :))V=9.83ワ!